OpenStudy (anonymous):

-6x + 8y = -60 -4x - 2y = 4

OpenStudy (anonymous):

@campbell_st Can you help? (:

OpenStudy (campbell_st):

multiply equation 2 by 4 and add the 2 equations to eliminate y - 6x + 8y = -60 -16x - 8y = 16 --------------- -22x = -44 divide by - 22 x = 2 substitute into either equation to find y -6 (2) + 8y = -60 8y = -48 y = -6

OpenStudy (anonymous):

Could you draw it on a graph? :p

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

just download Geogebra and you can use it to graph your equations http://geogebra.org/cms/ the best free graphing and geometry software available

OpenStudy (anonymous):

Thanks! (: